# Fusion of Multiscale Features Via Centralized Sparse-attention Network for EEG Decoding

### v1

@@ 2.4 Temporal Convolutional
The TCN structure we use follows the design of[20], with its core architecture shown in Figure 1.C. This module adopts a double residual structure, aiming to deeply extract high-level temporal features from brain electrical signals. The dilation factors for the two residual blocks are $d_{1}$ and $d_{2}$, respectively. Batch normalization and ELU activation functions are then applied to enhance training stability. Finally, after concatenating the data, it is fed into a fully connected layer for classification, yielding the final classification result:
@@ 3.1 BCIC-IV-2A
BCIC-IV-2A[43] provided by the Graz University of Technology, consists of recordings from nine healthy subjects. Each subject performed four MI tasks, including left hand, right hand, foot, and tongue movements. EEG signals were recorded using 22 Ag/AgCl electrodes at a sampling rate of 250 Hz. For each subject, two sessions were recorded on different days, serving as the training and testing sets, respectively. Each session comprises 288 trials, i.e., 72 trials per task. In our study, a 4-second time window was adopted for each trial[19].
@@ 3.2 BCIC-IV-2B
BCIC-IV-2B[44] also includes recordings from nine healthy subjects. Each subject performed two MI tasks, namely left-hand and right-hand movements. EEG signals were recorded using three bipolar electrodes (C3, Cz, and C4) at a sampling rate of 250 Hz. A 4-second time window was adopted for each trial[19]. Each subject completed five sessions: the first two sessions without feedback, each containing 120 trials, and the subsequent three sessions with feedback, each containing 160 trials. In our experiments, the first three sessions were used for training and the last two sessions for testing, as done in[24].
@@ 3.3 HGD
HGD[14] consists of recordings from 14 healthy subjects. Each subject performed four MI tasks, including left-hand, right-hand, both-feet movements, and rest (where a visual cue identical to that of the other tasks was presented on the screen, but subjects were instructed not to perform any movement). The original EEG signals were recorded using 128 channels at a sampling rate of 500 Hz, with a 4-second time window for each trial. Following[26], we selected 44 channels covering the motor cortex for EEG decoding, downsampled the signals to 250 Hz, and applied standard normalization. For each subject, the dataset includes both training and testing sessions: the training set comprises approximately 880 trials per subject, while the testing set contains about 160 trials per subject.
@@ 3.4 SEED
SEED[45] comprises emotion-related EEG signals from 15 subjects. In the experimental paradigm, 15 film clips were employed to induce three distinct emotional states, namely positive, neutral, and negative. For each subject, data acquisition was conducted across three sessions, with an interval of approximately one week between consecutive sessions. EEG signals were recorded from 62 electrodes at a sampling rate of 1000 Hz and subsequently downsampled to 200 Hz. The continuous signals were segmented into non-overlapping 1-second windows, resulting in 3394 trials per session. All EEG recordings were further processed using a band-pass filter in the range of 4–47 Hz. Consistent with the experimental protocol described in[19], five-fold cross-validation was performed within a session.
@@ 3.5 SEED-VIG
SEED-VIG[46] comprises driving fatigue related EEG signals from 23 subjects. Participants performed simulated driving tasks in a virtual-reality environment covering a variety of weather and road conditions, with a total recording duration of 118 minutes. EEG signals were acquired from 17 channels at an initial sampling rate of 1000 Hz and subsequently downsampled to 200 Hz. To suppress electronic interference, the recordings were band-pass filtered in the 1–50 Hz range and then normalized.
@@
Following[46], a threshold of 0.35 was used to distinguish between fatigued and alert states. Five-fold cross-validation was employed to partition the data into train and test sets.
@@ 3.6 Experimental Details
All datasets are implemented using Python 3.10 and the PyTorch framework, and experiments are conducted on an NVIDIA RTX 2080Ti GPU. The cross-entropy loss function and the Adam optimizer are used with a learning rate set to 0.0009. To ensure fair comparison with baseline models, the number of training epochs for BCIC-IV-2A, BCIC-IV-2B, and HGD is set to 2000, with a batch size of 64. For SEED and SEED-VIG, the batch size is set to 128[47], with training epochs of 1000 and 500, respectively. The parameter $S$ in S&R is set to 8. Following[19], no data augmentation is applied to SEED. Since each trial in SEED is segmented into 1s windows, the model’s pooling sizes ($P_{1},P_{2}$) are set to (4, 4), while other datasets follow the specifications in Table 2. To ensure reproducibility, a fixed random seed is used throughout all experiments.
@@ 4 Result
For the three MI-based datasets (BCIC-IV-2A, BCIC-IV-2B, and HGD), we compared our method with the baseline model EEGNet[15] as well as several SOTA models proposed between 2023 and 2025, including EEG-Conformer[19], ATCNet[20], ADFCNN[28], EEG TransNet[24], MSTFNet[23], EISATC-Fusion[48], MCMTNet[25], and TMSA-Net[26].
@@
As shown in Table 5, on HGD, the EEG-CSANet outperformed several SOTA approaches, achieving an average accuracy of 96.43%, which is 0.53% higher than that of the second-best method, TMSA-Net. The average Kappa coefficient reached 0.9542, exceeding TMSA-Net by 0.009. In addition, the low variance of 4.52 indicates that the proposed model maintains stable and consistent classification performance across all 14 subjects.
@@
On SEED, we compared EEG-CSANet with several representative methods evaluated under the same five-fold cross-validation protocol, including SVM[12], DGCNN[49], V-IAG[50], EEG-Conformer[19], and CU-GCN[51]. EEG-CSANet achieved an average accuracy of 96.03% and a Kappa coefficient of 94.04%, outperforming all competing methods.
On SEED-VIG, comparisons were conducted with EEGNet[15], CSF-GTNet[52], SFT-Net[53], EEG-Conformer[19], and GAT-CNN[47]. EEG-CSANet obtained an average accuracy of 90.56%, representing an improvement of 0.42% over the second-best method, GAT-CNN. Notably, due to pronounced inter-subject variability, fatigue-related samples in SEED-VIG inevitably exhibit severe class imbalance. Under this challenging condition, EEG-CSANet still achieved a Kappa coefficient of 0.7327, exceeding SFT-Net by 0.0157, which demonstrates its robustness and reliability in imbalanced fatigue-detection scenarios.
@@ 5.1 Ablation Experiments
BCIC-IV-2A 88.54 81.35 8.41 7.91 84.72 75.36
BCIC-IV-2B 91.09 89.02 8.48 7.03 82.18 78.04
HGD 96.52 93.75 3.01 5.37 95.42 91.67
SEED 95.92 96.03 2.18 1.94 94.01 94.04
SEED-VIG 90.53 89.93 6.37 6.76 73.27 71.53
@@ 5.3 Feature Visualization and Confusion Matrices
In order to further illustrate the efficacy of EEG-CSANet, the Uniform Manifold Approximation and Projection (UMAP) method is applied to visualize the classification features extracted from five datasets. Compared with t-SNE, UMAP offers higher computational efficiency and scalability for high-dimensional data, revealing clustering patterns and distribution characteristics in low-dimensional space[54]. As illustrated in Figure 5, prior to training, the extracted features are dispersed in the low-dimensional space, exhibiting substantial overlap between different classes and limited overall discriminative capability. Following model training, the feature distribution forms a more compact and orderly clustering structure, with markedly enhanced inter-class separability.
@@ 6.1 Baselines
EEGNet [15] is a lightweight convolutional neural network specifically tailored for EEG signal classification. By leveraging depthwise separable convolutions, it achieves an effective trade-off between performance and computational efficiency. Its architecture is simple, compact, and easy to implement or transfer across different tasks. Owing to these advantages, EEGNet has been widely adopted in various BCI applications, such as MI and Event-Related Potention decoding, and has become one of the benchmark models in deep learning for EEG analysis.
@@
EEG-Conformer [19] is a deep learning model specifically developed for EEG signal classification, which integrates CNNs with a Transformer architecture. In this framework, CNNs are employed to extract local spatiotemporal features, while the Transformer is utilized to model long-range temporal dependencies and global contextual information. By effectively combining the strengths of feature extraction and sequence modeling, EEG-Conformer provides a powerful and comprehensive representation of EEG data.
@@
ATCNet [20] is a compact and interpretable attention-based temporal convolutional network specifically designed for EEG motor imagery classification. It integrates principles of scientific machine learning by employing convolutional layers to extract spatiotemporal features, while multi-head self-attention mechanisms emphasize the most informative temporal segments. In addition, a TCN is incorporated to capture long-term temporal dependencies. With its concise architecture and strong interpretability, ATCNet demonstrates improved decoding performance for EEG signals.
@@
ADFCNN [28] is a deep learning model tailored for BCI. It adopts a dual-scale convolutional architecture to separately extract temporal features of EEG rhythms as well as global and fine-grained spatial features. To further enhance representation learning, a self-attention mechanism is introduced, enabling the model to dynamically weight and fuse multiscale information according to intrinsic feature similarities, thereby improving feature discriminability. This design effectively overcomes the limitations of traditional single-scale or traditional multiscale CNNs in feature extraction and fusion, offering stronger modeling capabilities for both spectral and spatial information.
@@
EEG-Transnet [24] is a EEG classification model that integrates CNNs with a self-attention mechanism. The model extracts multimodal temporal features from both the mean and variance dimensions, while a shared self-attention module captures global temporal dependencies. A convolutional encoder is then employed to fuse these features, thereby enhancing their discriminability. In addition, a signal S&R-based data augmentation strategy is introduced to further improve robustness and decoding performance.
@@
MSTFNet [23] is an end-to-end convolutional neural network proposed for motor imagery EEG classification. To overcome the limitations of traditional single-scale convolutions in feature extraction, MSTFNet is designed with four core modules: feature enhancement, multiscale temporal feature extraction, spatial feature extraction, and feature fusion. Through the collaborative extraction of multiscale spatiotemporal features and the adoption of refined fusion strategies, the model substantially improves the decoding of EEG signals.
@@
EISATC-Fusion [48] is a neural network model tailored to EEG decoding, which integrates Inception modules, multi-head self-attention, TCNs, and a layer fusion structure. Specifically, the model employs DS-Inception to extract multiscale frequency band features, incorporates cnnCosMSA to alleviate attention collapse and enhance interpretability, and applies depthwise separable convolutions to reduce parameter complexity. Furthermore, by combining both feature-level and decision-level fusion strategies, EISATC-Fusion achieves improved robustness in EEG signal decoding.
@@
MCMTNet [25] is a deep learning model capable of directly processing raw EEG signals without the need for complex preprocessing. Its architecture is composed of three main components: a multi-domain convolution module, a multi-head attention module, and a TCN. The multi-domain convolution module is responsible for filtering and feature extraction, the multi-head attention module enhances cross-scale correlations, and the TCN improves temporal coherence.
@@
TMSA-Net [26] is a neural network model that integrates CNNs with an improved Transformer-based attention mechanism. The model first employs CNNs to extract local spatiotemporal features, and then introduces a novel attention module to enhance global modeling capability across both the channel and temporal dimensions, thereby effectively bridging the gap between local and global representations. By optimizing the attention structure, TMSA-Net not only reduces computational overhead but also improves feature fusion efficiency, enhancing the model’s sensitivity to key EEG patterns and its interpretability.
@@
SVM[12] is a classical supervised learning model widely employed in small-sample, high-dimensional pattern recognition tasks such as EEG signal classification. Its core principle lies in finding an optimal hyperplane that maximizes the margin between samples of different classes, thereby yielding a decision boundary with strong generalization capability. In EEG analysis, SVM is typically combined with specific feature extraction methods such as time-frequency features or Common Spatial Patterns (CSP) and has demonstrated robust performance in EEG decoding tasks.
@@
DGCNN[49] is a dynamic graph convolutional neural network specifically designed for modeling EEG signals. Unlike traditional graph convolutional networks that rely on predefined or fixed adjacency matrices, DGCNN dynamically learns the connection weights between EEG channels during training, thereby adaptively constructing an adjacency matrix that captures the intrinsic functional relationships among channels. Thanks to its flexible graph structure learning capability and end-to-end trainability, DGCNN has demonstrated superior performance in various EEG analysis tasks such as emotion recognition and cognitive state decoding, and has become one of the key benchmark approaches for processing EEG signals with graph neural networks.
@@
V-IAG[50] is a graph-based modeling approach designed for EEG emotion recognition, aiming to jointly capture individual differences and dynamic uncertainties among brain regions. The method comprises two branches: an instance-adaptive branch that dynamically constructs a graph reflecting subject-specific channel dependencies based on the input, and a variational branch that generates a probabilistic graph to quantify connection uncertainty. The fused graph is then used to extract more discriminative features. To further enhance representation capability, V-IAG introduces a multi-level multi-graph convolution operation that aggregates channel features across different frequency bands using distinct graph structures, complemented by graph coarsening and sparsity constraints to improve robustness.
@@
CU-GCN[51] is a graph convolutional network for EEG emotion recognition that employs distribution-based uncertainty modeling to adaptively weight node features, effectively alleviating over-smoothing in long-range propagation while capturing spatial dependencies and temporal-spectral relationships among EEG channels. The model integrates graph mixup to enhance latent connections and suppress label noise, and incorporates the uncertainty mechanism into deep GCN weights via a one-way learning strategy, significantly outperforming all baseline models.
@@
CSF-GTNet[52] is a multi-dimensional feature fusion network designed for driver fatigue detection, which innovatively integrates time-domain and spatial–spectral information to enhance the representation capability of EEG signals. The model consists of two parallel branches: the Gaussian Time-domain Network (GTNet), which adaptively captures fatigue-related temporal dynamics, and the pure Convolutional Spatial–Frequency Network (CSFNet), which focuses on extracting multi-scale spatial distributions and spectral characteristics. By fusing features from these complementary dimensions, CSF-GTNet provides a more comprehensive characterization of the differences in EEG patterns between alert and fatigue states.
@@
SFT-Net[53] is a lightweight and efficient model for EEG-based driver fatigue detection that enhances recognition performance by fusing spatial, spectral, and temporal information. The method constructs a 4D feature tensor from the differential entropy of five frequency bands, employs an attention mechanism to recalibrate spatial and spectral features for each time frame, extracts key representations using depthwise separable convolution, and captures temporal dependencies via LSTM. SFT-Net outperforms existing methods and demonstrates strong interpretability, validating the effectiveness of multi-dimensional information fusion in fatigue detection.
@@
GAT-CNN[47] is an end-to-end EEG model for driver fatigue detection that enhances recognition performance by explicitly modeling dependencies among EEG channels. The method takes preprocessed EEG signals and an adjacency matrix constructed from mutual information as input, uses a graph attention mechanism to adaptively learn inter-channel weights, and combines convolutional operations to extract structured spatiotemporal features without manual feature engineering. Alert and fatigue states are finally classified using fully connected layers and a Softmax classifier. GAT-CNN demonstrates excellent predictive accuracy and generalization capability in experiments.

### v2

@@ 2.4 Temporal Convolutional
The TCN structure we use follows the design of[altaheri2022physics], with its core architecture shown in Figure 1.C. This module adopts a double residual structure, aiming to deeply extract high-level temporal features from brain electrical signals. The dilation factors for the two residual blocks are $d_{1}$ and $d_{2}$, respectively. Batch normalization and ELU activation functions are then applied to enhance training stability. Finally, after concatenating the data, it is fed into a fully connected layer for classification, yielding the final classification result:
@@ 3.1 BCIC-IV-2A
BCIC-IV-2A[brunner2008bci] provided by the Graz University of Technology, consists of recordings from nine healthy subjects. Each subject performed four MI tasks, including left hand, right hand, foot, and tongue movements. EEG signals were recorded using 22 Ag/AgCl electrodes at a sampling rate of 250 Hz. For each subject, two sessions were recorded on different days, serving as the training and testing sets, respectively. Each session comprises 288 trials, i.e., 72 trials per task. In our study, a 4-second time window was adopted for each trial[song2022eeg].
@@ 3.2 BCIC-IV-2B
BCIC-IV-2B[leeb2008bci] also includes recordings from nine healthy subjects. Each subject performed two MI tasks, namely left-hand and right-hand movements. EEG signals were recorded using three bipolar electrodes (C3, Cz, and C4) at a sampling rate of 250 Hz. A 4-second time window was adopted for each trial[song2022eeg]. Each subject completed five sessions: the first two sessions without feedback, each containing 120 trials, and the subsequent three sessions with feedback, each containing 160 trials. In our experiments, the first three sessions were used for training and the last two sessions for testing, as done in[ma2024attention].
@@ 3.3 HGD
HGD[schirrmeister2017deep] consists of recordings from 14 healthy subjects. Each subject performed four MI tasks, including left-hand, right-hand, both-feet movements, and rest (where a visual cue identical to that of the other tasks was presented on the screen, but subjects were instructed not to perform any movement). The original EEG signals were recorded using 128 channels at a sampling rate of 500 Hz, with a 4-second time window for each trial. Following[zhao2025tmsa], we selected 44 channels covering the motor cortex for EEG decoding, downsampled the signals to 250 Hz, and applied standard normalization. For each subject, the dataset includes both training and testing sessions: the training set comprises approximately 880 trials per subject, while the testing set contains about 160 trials per subject.
@@ 3.4 SEED
SEED[zheng2018emotionmeter] comprises emotion-related EEG signals from 15 subjects. In the experimental paradigm, 15 film clips were employed to induce three distinct emotional states, namely positive, neutral, and negative. For each subject, data acquisition was conducted across three sessions, with an interval of approximately one week between consecutive sessions. EEG signals were recorded from 62 electrodes at a sampling rate of 1000 Hz and subsequently downsampled to 200 Hz. The continuous signals were segmented into non-overlapping 1-second windows, resulting in 3394 trials per session. All EEG recordings were further processed using a band-pass filter in the range of 4-47 Hz. Consistent with the experimental protocol described in[song2022eeg], five-fold cross-validation was performed within a session.
@@ 3.5 SEED-VIG
SEED-VIG[zheng2017multimodal] comprises driving fatigue related EEG signals from 23 subjects. Participants performed simulated driving tasks in a virtual-reality environment covering a variety of weather and road conditions, with a total recording duration of 118 minutes. EEG signals were acquired from 17 channels at an initial sampling rate of 1000 Hz and subsequently downsampled to 200 Hz. To suppress electronic interference, the recordings were band-pass filtered in the 1-50 Hz range and then normalized.
@@
Following[zheng2017multimodal], a threshold of 0.35 was used to distinguish between fatigued and alert states. Five-fold cross-validation was employed to partition the data into train and test sets.
@@ 3.6 Experimental Details
All datasets are implemented using Python 3.10 and the PyTorch framework, and experiments are conducted on an NVIDIA RTX 2080Ti GPU. The cross-entropy loss function and the Adam optimizer are used with a learning rate set to 0.0009. To ensure fair comparison with baseline models, the number of training epochs for BCIC-IV-2A, BCIC-IV-2B, and HGD is set to 2000, with a batch size of 64. For SEED and SEED-VIG, the batch size is set to 128[chen2025driver], with training epochs of 1000 and 500, respectively. The parameter $S$ in S&R is set to 8. Following[song2022eeg], no data augmentation is applied to SEED. Since each trial in SEED is segmented into 1s windows, the model’s pooling sizes ($P_{1},P_{2}$) are set to (4, 4), while other datasets follow the specifications in Table 2. To ensure reproducibility, a fixed random seed is used throughout all experiments.
@@ 4 Result
For the three MI-based datasets (BCIC-IV-2A, BCIC-IV-2B, and HGD), we compared our method with the baseline model EEGNet[lawhern2018eegnet] as well as several SOTA models proposed between 2023 and 2025, including EEG-Conformer[song2022eeg], ATCNet[altaheri2022physics], ADFCNN[tao2023adfcnn], EEG TransNet[ma2024attention], MSTFNet[jin2024multiscale], EISATC-Fusion[liang2024eisatc], MCMTNet[yang2025mcmtnet], and TMSA-Net[zhao2025tmsa].
@@
As shown in Table 5, on HGD, the EEG-CSANet outperformed several SOTA approaches, achieving an average accuracy of 97.15%, which is 1.25% higher than that of the second-best method, TMSA-Net. The average Kappa coefficient reached 0.9627, exceeding TMSA-Net by 0.018. In addition, the low STD of 2.97 indicates that the proposed model maintains stable and consistent classification performance across all 14 subjects.
@@
On SEED, we compared EEG-CSANet with several representative methods evaluated under the same five-fold cross-validation protocol, including SVM[zheng2015investigating], DGCNN[song2018eeg], V-IAG[song2021variational], EEG-Conformer[song2022eeg], and CU-GCN[gao2024graph]. EEG-CSANet achieved an average accuracy of 96.03% and a Kappa coefficient of 94.04%, outperforming all competing methods.
On SEED-VIG, comparisons were conducted with EEGNet[lawhern2018eegnet], CSF-GTNet[gao2023csf], SFT-Net[gao2023sft], EEG-Conformer[song2022eeg], and GAT-CNN[chen2025driver]. EEG-CSANet obtained an average accuracy of 90.56%, representing an improvement of 0.42% over the second-best method, GAT-CNN. Notably, due to pronounced inter-subject variability, fatigue-related samples in SEED-VIG inevitably exhibit severe class imbalance. Under this challenging condition, EEG-CSANet still achieved a Kappa coefficient of 0.7327, exceeding SFT-Net by 0.0157, which demonstrates its robustness and reliability in imbalanced fatigue-detection scenarios.
@@ 5.1 Ablation Experiments
BCIC-IV-2A 88.54 81.35 8.41 7.91 0.8472 0.7536
BCIC-IV-2B 91.09 89.02 8.48 7.03 0.8218 0.7804
HGD 97.15 93.75 2.97 5.37 0.9627 0.9167
SEED 95.92 96.03 2.18 1.94 0.9401 0.9404
SEED-VIG 90.53 89.93 6.37 6.76 0.7327 0.7153
@@ 5.3 Feature Visualization and Confusion Matrices
In order to further illustrate the efficacy of EEG-CSANet, the Uniform Manifold Approximation and Projection (UMAP) method is applied to visualize the classification features extracted from five datasets. Compared with t-SNE, UMAP offers higher computational efficiency and scalability for high-dimensional data, revealing clustering patterns and distribution characteristics in low-dimensional space[mcinnes2018umap]. As illustrated in Figure 5, prior to training, the extracted features are dispersed in the low-dimensional space, exhibiting substantial overlap between different classes and limited overall discriminative capability. Following model training, the feature distribution forms a more compact and orderly clustering structure, with markedly enhanced inter-class separability.
@@ 6.1 Baselines
EEGNet [lawhern2018eegnet] is a lightweight convolutional neural network specifically tailored for EEG signal classification. By leveraging depthwise separable convolutions, it achieves an effective trade-off between performance and computational efficiency. Its architecture is simple, compact, and easy to implement or transfer across different tasks. Owing to these advantages, EEGNet has been widely adopted in various BCI applications, such as MI and Event-Related Potention decoding, and has become one of the benchmark models in deep learning for EEG analysis.
@@
EEG-Conformer [song2022eeg] is a deep learning model specifically developed for EEG signal classification, which integrates CNNs with a Transformer architecture. In this framework, CNNs are employed to extract local spatiotemporal features, while the Transformer is utilized to model long-range temporal dependencies and global contextual information. By effectively combining the strengths of feature extraction and sequence modeling, EEG-Conformer provides a powerful and comprehensive representation of EEG data.
@@
ATCNet [altaheri2022physics] is a compact and interpretable attention-based temporal convolutional network specifically designed for EEG motor imagery classification. It integrates principles of scientific machine learning by employing convolutional layers to extract spatiotemporal features, while multi-head self-attention mechanisms emphasize the most informative temporal segments. In addition, a TCN is incorporated to capture long-term temporal dependencies. With its concise architecture and strong interpretability, ATCNet demonstrates improved decoding performance for EEG signals.
@@
ADFCNN [tao2023adfcnn] is a deep learning model tailored for BCI. It adopts a dual-scale convolutional architecture to separately extract temporal features of EEG rhythms as well as global and fine-grained spatial features. To further enhance representation learning, a self-attention mechanism is introduced, enabling the model to dynamically weight and fuse multiscale information according to intrinsic feature similarities, thereby improving feature discriminability. This design effectively overcomes the limitations of traditional single-scale or traditional multiscale CNNs in feature extraction and fusion, offering stronger modeling capabilities for both spectral and spatial information.
@@
EEG-Transnet [ma2024attention] is a EEG classification model that integrates CNNs with a self-attention mechanism. The model extracts multimodal temporal features from both the mean and variance dimensions, while a shared self-attention module captures global temporal dependencies. A convolutional encoder is then employed to fuse these features, thereby enhancing their discriminability. In addition, a signal S&R-based data augmentation strategy is introduced to further improve robustness and decoding performance.
@@
MSTFNet [jin2024multiscale] is an end-to-end convolutional neural network proposed for motor imagery EEG classification. To overcome the limitations of traditional single-scale convolutions in feature extraction, MSTFNet is designed with four core modules: feature enhancement, multiscale temporal feature extraction, spatial feature extraction, and feature fusion. Through the collaborative extraction of multiscale spatiotemporal features and the adoption of refined fusion strategies, the model substantially improves the decoding of EEG signals.
@@
EISATC-Fusion [liang2024eisatc] is a neural network model tailored to EEG decoding, which integrates Inception modules, multi-head self-attention, TCNs, and a layer fusion structure. Specifically, the model employs DS-Inception to extract multiscale frequency band features, incorporates cnnCosMSA to alleviate attention collapse and enhance interpretability, and applies depthwise separable convolutions to reduce parameter complexity. Furthermore, by combining both feature-level and decision-level fusion strategies, EISATC-Fusion achieves improved robustness in EEG signal decoding.
@@
MCMTNet [yang2025mcmtnet] is a deep learning model capable of directly processing raw EEG signals without the need for complex preprocessing. Its architecture is composed of three main components: a multi-domain convolution module, a multi-head attention module, and a TCN. The multi-domain convolution module is responsible for filtering and feature extraction, the multi-head attention module enhances cross-scale correlations, and the TCN improves temporal coherence.
@@
TMSA-Net [zhao2025tmsa] is a neural network model that integrates CNNs with an improved Transformer-based attention mechanism. The model first employs CNNs to extract local spatiotemporal features, and then introduces a novel attention module to enhance global modeling capability across both the channel and temporal dimensions, thereby effectively bridging the gap between local and global representations. By optimizing the attention structure, TMSA-Net not only reduces computational overhead but also improves feature fusion efficiency, enhancing the model’s sensitivity to key EEG patterns and its interpretability.
@@
SVM[zheng2015investigating] is a classical supervised learning model widely employed in small-sample, high-dimensional pattern recognition tasks such as EEG signal classification. Its core principle lies in finding an optimal hyperplane that maximizes the margin between samples of different classes, thereby yielding a decision boundary with strong generalization capability. In EEG analysis, SVM is typically combined with specific feature extraction methods such as time-frequency features or Common Spatial Patterns (CSP) and has demonstrated robust performance in EEG decoding tasks.
@@
DGCNN[song2018eeg] is a dynamic graph convolutional neural network specifically designed for modeling EEG signals. Unlike traditional graph convolutional networks that rely on predefined or fixed adjacency matrices, DGCNN dynamically learns the connection weights between EEG channels during training, thereby adaptively constructing an adjacency matrix that captures the intrinsic functional relationships among channels. Thanks to its flexible graph structure learning capability and end-to-end trainability, DGCNN has demonstrated superior performance in various EEG analysis tasks such as emotion recognition and cognitive state decoding, and has become one of the key benchmark approaches for processing EEG signals with graph neural networks.
@@
V-IAG[song2021variational] is a graph-based modeling approach designed for EEG emotion recognition, aiming to jointly capture individual differences and dynamic uncertainties among brain regions. The method comprises two branches: an instance-adaptive branch that dynamically constructs a graph reflecting subject-specific channel dependencies based on the input, and a variational branch that generates a probabilistic graph to quantify connection uncertainty. The fused graph is then used to extract more discriminative features. To further enhance representation capability, V-IAG introduces a multi-level multi-graph convolution operation that aggregates channel features across different frequency bands using distinct graph structures, complemented by graph coarsening and sparsity constraints to improve robustness.
@@
CU-GCN[gao2024graph] is a graph convolutional network for EEG emotion recognition that employs distribution-based uncertainty modeling to adaptively weight node features, effectively alleviating over-smoothing in long-range propagation while capturing spatial dependencies and temporal-spectral relationships among EEG channels. The model integrates graph mixup to enhance latent connections and suppress label noise, and incorporates the uncertainty mechanism into deep GCN weights via a one-way learning strategy, significantly outperforming all baseline models.
@@
CSF-GTNet[gao2023csf] is a multi-dimensional feature fusion network designed for driver fatigue detection, which innovatively integrates time-domain and spatial-spectral information to enhance the representation capability of EEG signals. The model consists of two parallel branches: the Gaussian Time-domain Network (GTNet), which adaptively captures fatigue-related temporal dynamics, and the pure Convolutional Spatial-Frequency Network (CSFNet), which focuses on extracting multi-scale spatial distributions and spectral characteristics. By fusing features from these complementary dimensions, CSF-GTNet provides a more comprehensive characterization of the differences in EEG patterns between alert and fatigue states.
@@
SFT-Net[gao2023sft] is a lightweight and efficient model for EEG-based driver fatigue detection that enhances recognition performance by fusing spatial, spectral, and temporal information. The method constructs a 4D feature tensor from the differential entropy of five frequency bands, employs an attention mechanism to recalibrate spatial and spectral features for each time frame, extracts key representations using depthwise separable convolution, and captures temporal dependencies via LSTM. SFT-Net outperforms existing methods and demonstrates strong interpretability, validating the effectiveness of multi-dimensional information fusion in fatigue detection.
@@
GAT-CNN[chen2025driver] is an end-to-end EEG model for driver fatigue detection that enhances recognition performance by explicitly modeling dependencies among EEG channels. The method takes preprocessed EEG signals and an adjacency matrix constructed from mutual information as input, uses a graph attention mechanism to adaptively learn inter-channel weights, and combines convolutional operations to extract structured spatiotemporal features without manual feature engineering. Alert and fatigue states are finally classified using fully connected layers and a Softmax classifier. GAT-CNN demonstrates excellent predictive accuracy and generalization capability in experiments.